OpenStudy (anonymous):

slope intercept form of equation of line through points. through: (-5, -3), slope = 2/5

myininaya (myininaya):

equation of a line can be written in the form (x-x_1)m=y_1-y

myininaya (myininaya):

oops (x-x_1)m=y-y_1

myininaya (myininaya):

we know m=2/5 and we know a point (x_1,y_1)=(-5,-3)

myininaya (myininaya):

(x-(-5))2/5=(y-(-3))

myininaya (myininaya):

(x+5)2/5=y+3

myininaya (myininaya):

2x/5+5*2/5=y+3

myininaya (myininaya):

2x/5+2=y+3

myininaya (myininaya):

subtract 3 on both sides to put it in the form y=mx+b

OpenStudy (anonymous):

okay

myininaya (myininaya):

y=2x/5-1

OpenStudy (anonymous):

i just dont understand how -5 and -3 turned into -1

myininaya (myininaya):

if we subtract 3 on both sides 2-3=-1

OpenStudy (anonymous):

of mx + b?

myininaya (myininaya):

no you were asking how i got -1 right?

OpenStudy (anonymous):

yea i understand the y= 2/5 x just not how u got -1 from (-5, -3)

myininaya (myininaya):

do you understand where I got 2x/5+2=y+3

OpenStudy (anonymous):

not the 2nd 2 and 3

myininaya (myininaya):

the point slope formula is (x-x_1)m=y-y_1 where in this case a line on this point is (x_1,y_1)=(-5,-3) and m=2/5 since that is the slope of this line

myininaya (myininaya):

so we have (x-(-5))*2/5=y-(-3)

myininaya (myininaya):

(x+5)*2/5=y+3

myininaya (myininaya):

Use distributive property we have 2x/5+5*2/5=y+3

myininaya (myininaya):

we want to put this in y=mx+b form which is the slope intercept form

myininaya (myininaya):

Since the above equation has y by itself we need to solve for y by subtracting 3 on both sides

myininaya (myininaya):

2x/5+2-3=y

myininaya (myininaya):

y=2x/5+2-3

myininaya (myininaya):

y=2x/5-1

OpenStudy (anonymous):

o okay now i understand thank u

myininaya (myininaya):

cool!

OpenStudy (amistre64):

the slope-intercept equation is: y = mx + b You know the slope, and they give you (x=-5,y=-3); so lets plug and play... -3 = (2/5)(-5) +b ; solve for "b". -3 = -2 + b -1 = b throw away your (-5,-3) and use the equation as it: y = (2/5)x - 1